Plan for E/D Night UT 03-Mar-2020

  • Engineering Tasks
    • SOUL SX AO
    • General Telescope
      • PCS patch 20200226 for TCS 2020B (needs to be activated by OSA)
        • If any problems, roll back patch - no troubleshooting
        • PCS patch 20200226 for TCS 2020B release is available (but not yet activated). This patch hopefully fixes the problem of co-pointing error after instrument authorization changes (see IT#7380). To activate the patch, do the following in a command-line window:

          1. ssh tcs@tcs1.mountain.lbto.org
          2. cd /lbt/tcs/current
          3. ./tcs-build latest PCS (it is OK to answer "n" to the restart question. This way, patch will be run after TO restart affected subsystem.)

          In case of any problem, please revert it back. To revert the patch, enter the following commands:

          1. ssh tcs@tcs1.mountain.lbto.org
          2. cd /lbt/tcs/current
          3. ./tcs-build revert PCS
      • Confirmation on-sky AGW-2 WFS spots are gone - 10min
    • LBC Testing
      • Mixed-mode LBC-R/LUCI-1 tests - dithering tests - medium priority - 0.5 hr - presets worked in full binocular mode - can be done in poor weather
      • Photometric standards - requires clear skies but any seeing
    • LUCI-1 (LUCI-2 off telescope)
      • Photometric standards - requires clear skies but any seeing
    • MODS
      • Photometric standards - requires clear skies but any seeing
  • Science Targets
    • Q1237+134 - MODS Bino - 1hr (O. Kuhn)
      • Files located: OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/
    • Quasars @RA 10hr and 12hr - MODS -Bino -1hr (O. Kuhn)
    • Novae Targets - MODS-Bino Dual Grating - 0.75 hours per target (M. Wagner)
      • Files located: OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/
        • V392 Per 04:43:21.37 47:21:25.87
          V2860 Ori 06:09:57.453 12:12:25.15
          V959 Mon 06:39:38.61 05:49:26.74
          V435 CMa 07:13:45.84 -21:12:31.3
    • Triple Merger Candidates - LUCI-1 (B. Rothberg)
      • Files located: OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/Triple_AGNs_Rothberg_LD_2019B_007
      • LUCI observations have priority over MODS
      • Triple Merger Candidate SDSSJ 0838+5055 - visible twilight untl 2am (hits 30deg limit)
        • LUCI-1 - 2.75 hrs G200 1" slit HKspec
        • MODS-Bino - 1 hr Dual Grating 1" slit
        • Requires 1" or better on AGW/WFS
      • Triple Merger Candidate SDSSJ 0853+5245- visible twilight untl 2am (hits 30deg limit)
        • LUCI-1 - 2.75 hrs G200 1" slit HKspec
        • MODS-Bino - 1 hr Dual Grating 1" slit
        • Requires 1" or better on AGW/WFS
      • Triple Merger Candidate SDSS J1342+19133- all 2nd half - higher priority
        • LUCI-1 - 2.75 hrs G200 1" slit HKspec
        • MODS-Bino - 1 hr Dual Grating 1" slit
        • nuclei separations are 5" and 8" - so looser constraints
    • NGC 2403-5 - MODS-Bino -1.5-2.5 hr some clouds seeing < 1.5" (M. Edwards)
      • Files located: OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/
  • Poor Weather Backup Targets
    • B. Rothberg
      • Local Mergers - MODS Bino or LUCI-2
        • Files are located on OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/Rothberg_Local_Mergers/
        • OR RM507-5/4: /home/observer/share/Service_Observing/Eng_Jan_2020/Rothberg_Local_Mergers/
        • MODS (dual grating)
          • UGC 4079 - 07:55:06 +55:42:13 - 1 hour
          • NGC 3310 - 10:38:45 +53:30:11 - 1 hour
          • NGC 2744 - 09:04:39 +18:27:52 - 1 hour
          • NGC 6598 - 18:08:56 +69:04:04 - 1 hour (end of night)
        • LUCI-1
          • Files are located on OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/Rothberg_Local_Mergers/
          • OR on RM507-5: /home/observer/share/Service_Observing/Eng_Jan_2020/Rothberg_Local_Mergers/
          • Arp 187 - 05:04:53 -10:14:52 - visible until 11:30pm - 2 hours (covers zJ and HKspec observations) - G200 HK/zJspec observations
          • Arp 156 - 10:42:38 +77:29:41 (requires the CWL 2.3 cals) G210 K-band
      • ICRF3 project with USNO - LUCI-1
        • targets available all night
      • H-band Library (poorest conditions) - LUCI-1
        • STAR: RA Dec Class Hmag
        • HD 20797 03 24 40.55 +64 35 09.57 M0Iab 0.5
        • HD 42543 06 12 19.10 +22 54 30.7 M2I 0.9
        • HD 44537 06 24 53.90 +49 17 16.4 M2I 0.8
        • BD -15 1319 06 15 01.88 -15 52 49.7 K5II 6.4
  • Closed Dome Tasks
    • DONE LUCI-1 checks - 0.5-1hr (D. Thompson)
    • SX SOUL NCPA - 9pm (local) onwards (D. Miller, S. Walsh) - Requires retro reflector
      • DONE RetroReflector (RR) to be installed IF night is called or before 8pm
      • If RR not availble - use ARGOS cal unit - but susceptible to vibrations

-- BarryRothberg - 02 Mar 2020
Topic revision: r4 - 03 Mar 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback